Sci IPT 15 April 2009

See Unredacted notes for more complete information

Sci IPT 15 April 2009 Hills, Morita, Paola, Mangum, Emerson C Wilson Nyman Corder Brogan Remijan Testi

Photo of the AEM antenna shown there. Await mount, due imminently. Second AEM BUS arrives at the site (yesterday afternoon). WVR unconditionally accepted.

Power supply negotiations ran through the last year to establish power line to the Chilean grid at Calama. Those negotiations have fallen through. ALMA is now focussed on a gas turbine generator. LPG would be the fuel of choice. This is the power source for the VLT. Move to alternative power sources or wind energy is still being investigated. This also means a fiber will not be run along the power line.

Review of commissioning plans. A meeting had been planned for 18-19 June now but could be later. RH suggests 6-10 July. Neither Testi nor Wootten can attend at that time.

Recent work on site. Fringes achieved on the beacon. Too many obstacles to achieve astronomical fringes at the moment. Progress is quite slow. New problems crop up faster than existing problems are being fixed, still ascending the curve looking forward to rounding the top. As the little engine said, 'We think we can, we think we can." Routine total power observations of planets. OPT is working less reliably. Software has many new features to cope with new equipment, getting this working occupies quite a bit of time.

Mangum. I think there is a problem with PM03 aside from the OPT. RH: There are problems with the servo at the moment. However, the camera is also sick. Mangum: Saito and Ohishi are working on this problem. RH: We anticipate their discoveries with interest.

Change requests in FE area. One is providing IR filters and windows with a tilt with respect to the beams to cut down on reflections. Although designed to had low reflections, standing waves are seen. B3 optics consist of flat and curved mirror. At one end of the band, efficiency is low owing to beam being too narrow. Proposal is to curve both mirrors to raise efficiency from 80 to 85%. Should this be pushed to the point where we get the maximum aperture efficiency. This pushes things to a lower beam efficiency. So how much of the time will we observe point sources. Some discussion on this point. Suggest examine DRSP as one element of input.

New draft of science requirements is being worked on between Tom and Al. New requirement on operational efficiency. We have referred to this as latency. For science specs an appropriate phrase would be that latency should be such that loss of S/N should be negligible. Then the discussion would be over what the value of negligible actually is but that would presumably be defined outside the controlled document. DTE--we cannot change fast slew time of the antennas. One approach would be to say software latency should not undercut hardware latency. BD: Also the issue of calibration time. RH: That depends on getting the techniques right. But the time for a calibration does involve the software doing many different things. DTE: Additional loss of time should not dominate the mechanical time to get to source. The mechanical movement should be the limit. RH: We should budget hardware delays then say software should be a small fraction of this. RL: There was a diagram of this at the last system review. That did not make it into software requirements. RH: If anyone can find this document please bring it to light. RM: Is there a standard on the weather decision timescale? RH: I think that is 30 minutes. There was a spec of 2s on the cal wheel which was changed without process to 9s. Tests suggest 7s.

Calibration. Mangum working on new weather station, now at OSF. ALMA Cal Devices (ACD)s cleared PAI. These have the interim hot load on them. No word on a definitive production design for hot load.

WVRs--one at OSF, one in Garching in the lab. One in Garching having software interfaces tested, then goes to Onsala and to OSF. Will try to do atmospheric dips from lab window. Measured about 5mm PWV out the window, which is what is expected at the OSF during April (ref http://tinyurl.com/c3banq)

eSSR. Requirements review will be held 1-2 June. Telecon discussion next week. f2f meeting in Garching two weeks ago was held.

DRSP table in review. LT sent Copy to AW.

RH has a student who could look at site testing phase data. Oxygen sounder is moving ahead. LT: Agreement signed. Kickoff meeting with John and Boyan 30 April, the activity will begin then. A contract to be developed to give tech spec for unit being bought, to develop data acquisition software and to deploy on Chajnantor. Weather stations are on site, first installed at OSF. LN: even if software not up, data to archive. RH: Good. ICD defines data link.

Remijan reported on Splatalogue v2 upgrades; see agenda.

White Papers on science-driven instrumental development for the AS2010 Decadal Review were due on the weekend past. Some NRAO coauthored contributions, some involving ALMA, may be found at the agenda link. There were also instrumental white papers. Ones with potential ALMA interest include those on high frequency receiver development, array receiver development and improved atmospheric phase correction. Please write AW for a copy of the latter, which was prepared by many groups under CARMA leadership.

Next Sci IPT telecon 20 May.

-- AlWootten - 2009-04-15
Topic revision: r1 - 2009-04-15, AlWootten
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback